# Bounds on the charge of the graviton using gravitational wave observations

### v1

@@ III Detector dependent phase shifts due to graviton propagation
$\Delta\phi=\frac{q_{g}\Phi}{\hbar c}\leavevmode\nobreak\ .$ (12)
Where $\Phi$ is the total flux in the region enclosed by the two trajectories. For randomly oriented magnetic fields with a typical magnitude $B$ and correlation length $\lambda_{B}$, we have $D_{L}/\lambda_{B}$ magnetic field domains between the source located at a distance $D_{L}$ away and the observer. Since the magnetic field domains are randomly oriented, only one-third of all the domains will contribute to the phase, behaving like a random walk. We can compute the mean distance from the origin for this random walk after $D_{L}/3\lambda_{B}$ steps to be $\sqrt{2D_{L}/3\pi\lambda_{B}}$. Then, the total flux in the triangular region as illustrated in Fig. (2) reads Altschul (2007, 2008)
$\Phi=\sqrt{\frac{D_{L}\lambda_{B}}{6\pi}}Bd\leavevmode\nobreak\ ,\leavevmode\nobreak\ q_{g}=\dfrac{\sqrt{6\pi}\hbar c}{\sqrt{\lambda_{B}}B}\quantity(\dfrac{\Delta\phi}{\sqrt{D_{L}}d})$ (13)
@@
Here, $d$ is the separation between two detectors. The phase difference derived above is associated with distinct but coherent gravitons generated at the source observed at two different detectors. Nevertheless, similar to the case of photons, we can equate the classical GW phase difference observed and quantum phases of a graviton through the correspondence principle Altschul (2007, 2008). So, by measuring the phase difference between the GW perturbations received at two detectors and associating this phase shift with the quantum phase shift of the graviton, we should be able to place bounds on the charge of the graviton using the above equation. Note that unlike the prescription described in the previous section, these bounds are free of any degeneracies with the mass of the graviton.
@@
In general, measuring $\Delta\phi$ between two detectors for a particular GW signal is difficult. For equal-mass, face-on GW signals that are well-described by the dominant $(\ell=m=2)$ harmonic of the GW signal, $\Delta\phi$ is perfectly degenerate with the reference phase at coalescence $\phi_{c}$, and any finite $\Delta\phi$ can be absorbed as a redefinition of $\phi_{c}$. However, when non-quadropolar modes are relevant for describing the GW signal (e.g. for unequal mass binaries or highly inclined binaries), the degeneracy between $\phi_{c}$ and $\Delta\phi$ is broken, and it is possible to measure $\Delta\phi$. This effect is already well-known in the literature in works dealing with strong gravitational lensing of GWs. Strong lensing naturally produces “Type-II” images that have an a phase shift of $\pi/2$ relative to an un-lensed signal Schneider et al. (1992); Dai and Venumadhav (2017), and this phase shift has been shown to be measurable in presence of higher harmonics of the GW radiation, orbital precession and eccentricity in the GW signal Ezquiaga et al. (2021); Wang et al. (2021); Vijaykumar et al. (2023); Janquart et al. (2021). Hence, by making appropriate choices of the GW source, we can bound $\Delta\phi$ and thus the graviton charge independent of the mass.
IV Constraining the graviton charge
IV.1 Assumptions on $B$ and $\lambda_{B}$
@@
It is apparent from Eq. (11) and Eq. (13) that the bounds on $q_{g}$ will become tighter with increasing $B$ and $\lambda_{B}$. While the value of $B$ and $\lambda_{B}$ is not known, many observations have placed lower limits on these parameters (see Neronov et al. (2013); Subramanian (2016) for a review). For instance, Ref. Neronov and Vovk (2010) obtained a lower limit on $B$ exploiting the non-observation of GeV gamma-ray emission from the electromagnetic cascade initiated by TeV gamma-rays (from blazars) to be $B>3\times 10^{-16}\,{\rm G}$ while being coherent over Mpc scales Neronov and Vovk (2010). More specifically, these constraints state,
$B>\begin{cases}3\times 10^{-16}\quantity(\dfrac{\lambda_{B}}{1\,{\rm Mpc}})^{-1/2}\,{\rm G}\,,&\lambda_{B}\leq 1\,{\rm Mpc}\\ 3\times 10^{-16}\,{\rm G}\,,&\lambda_{B}>1\,{\rm Mpc}\end{cases}$ (14)
@@
While there have been some improvements to these constraints over the years, for our analysis we choose we shall make a conservative choice $B=3\times 10^{-16}\,{\rm G}$ and $\lambda_{B}=1\,{\rm Mpc}$ for our results.
IV.2 Results
Refer to captionRefer to caption
@@ IV.2 Results
Figure 3: (Top) Joint posterior on $m_{g}$ and $q_{g}/e$ parameter space as inferred from GWTC-3 data. The colour corresponds to the probability density—darker colour signifies a higher probability density. Since the measurement of $q_{g}$ and $m_{g}$ are degenerate following Eq. (10), we see that the isoprobability contours are ellipses in this two-dimensional $m_{g}$$q_{g}/e$ plane. The black dashed line is the 90% isoprobability contour. Marginalizing over $m_{g}$, we infer a 90% upper limit of $|q_{g}|/e<3\times 10^{-34}{}$. (Bottom) Posterior probability distribution on $q_{g}/e$ as estimated from the measured phase difference between LIGO-Hanford and LIGO-Livingston for GW190814. At 90% CL, we place an upper limit of $2\times 10^{-26}{}$ on $q_{g}/e$. Although weaker than the constraints obtained from reinterpreting the graviton mass constraints, these constraints are independent and are not degenerate with constraints on the graviton mass.
@@
We use the posterior on the mass of the (uncharged) graviton combined over GWTC-3 events by LVK collaboration Abbott et al. (2021c) and reinterpret it in the context of a charged massive graviton as discussed above. We report constraints on the joint $m_{g}$$q_{g}$ parameter space, with results shown in Fig. 3. As we see, the isoprobability contours in the image form ellipses, which follows directly from Eq. (11). We also show the ellipse corresponding to the $90\%$ contour. Under the aforementioned assumptions on $B$ and $\lambda_{B}$, the charge of the graviton is constrained to be
|$\absolutevalue{q_{g}}/e<3\times 10^{-34}{}\mbox{\quad.\quad}$ (15)
@@
We will next constrain the charge independently of the mass using Eq. (13). This method obviously has the advantage that we make no reference to the constraints on the graviton mass since the Aharonov-Bohm type phase shift can only occur in the presence of a non-zero charge. Even though this method is independent, the constraints are in general weaker than the one obtained using GW dephasing. As mentioned earlier, the measurement of phase difference between GW detectors can only be achieved with events that have significant higher harmonic content, failing which the phase difference measurement is exactly degenerate with the measurement of the coalescence phase of the binary $\phi_{c}$. We hence place bounds on $q_{g}$ using GW190814 Abbott et al. (2020), an asymmetric mass binary that showed significant evidence for higher-mode content in the signal. Although GW190814 was detected as a coincident signal by LIGO-Hanford, LIGO-Livingston, and Virgo detectors, we only use data from the LIGO detectors for our constraint due to the relatively low SNR in the Virgo detector222 While we have used only the phase difference between two detectors for our constraints with GW190814, this prescription can be easily generalized to multiple detectors by parametrizing all the relative phase differences through $q_{g}$. . We estimate the posterior on $\Delta\phi$ using the bilby Ashton et al. (2019) software package, using the nested sampler dynesty Speagle (2020) to sample the posterior distribution. We use the waveform approximant IMRPhenomXPHM Pratten et al. (2021) implemented in lalsuite LIGO Scientific Collaboration et al. (2018), and use the same standard priors as used in Ref. Abbott et al. (2020) along with a flat prior on $\Delta\phi$ in the range $[-\pi,\pi]$. We use the relative binning likelihood Cornish (2010); Zackay et al. (2018); Cornish (2021) implemented in bilby Krishna et al. (2023) to speed up the parameter inference [^3] While the relative binning implementation in bilby is strictly optimal for signals dominated by the quadrupolar mode of GW radiation, we have verified that the likelihood errors introduced by the non-optimality are small for our purposes. This is also in agreement with results of Ref. Krishna et al. (2023)., and use bilby_pipe Romero-Shaw et al. (2020) to streamline the inference runs. We obtain $\Delta\phi=-0.05^{+0.2}_{-0.25}$ at 90% CL which we consequently use to place an bound
|$\absolutevalue{q_{g}}/e<2\times 10^{-26}{}$ (16)
@@ V Conclusion
In this paper, we considered the possibility of a graviton endowed with a non-zero $U(1)$ charge. In the light of the Weinberg-Witten theorem, a non-zero $U(1)$ charge may make sense if the graviton is massive. A non-zero magnetic field will affect the trajectory of a charged graviton; we use this fact to compute the correction to the GW phasing that arises due to a non-zero charge. Our analysis points at a degeneracy of the graviton mass bounds from Abbott et al. (2021c) with the possible graviton charge. Using the LVK bounds on the mass of the graviton, we can impose an upper bound on the U(1) charge of a graviton in terms of the electron charge as $|q_{g}|/e<3\times 10^{-34}{}$.
@@
In addition to the above bound, we also use the total phase difference between the GW detections at LIGO-Hanford and LIGO-Livingston for the event GW190814 to find another independent upper bound of the graviton charge. This is done by identifying the phase difference with the possible Aharonov-Bohm phase shift experienced by a $U(1)$ charged graviton Altschul (2008, 2007). This provides us a novel bound on the graviton charge independent of its mass as $|q_{g}|/e<2\times 10^{-26}{}$. Similar to the bound obtained on the charge of a photon, the derivation of this independent bound on the graviton charge assumes the quantum phase difference arising due to a non-zero charge coincides with the classical phase difference of the two separate GW detections through the correspondence principle.
@@
In the near future, the constraints on $q_{g}$ will improve thanks to the larger baseline $d$ afforded by the addition of LIGO-India Unnikrishnan (2013); Saleem et al. (2022) to the GW detector network. These constraints will also improve by a few orders of magnitude with even larger baselines and/or sensitivities afforded by the next generation of ground-based (Cosmic Explorer Reitze et al. (2019), Einstein Telescope Punturo et al. (2010), etc.) and space-based (LISA Danzmann and Rudiger (2003), DECIGO Sato et al. (2017), etc.) detectors. Additionally, detections of more binaries with significant higher harmonics, eccentricity, and orbital precession will help in the measurement of the relative phase difference between detectors, hence also providing better constraints on $q_{g}$. Our analysis uses conservative estimates for the intergalactic magnetic field strength and the correlation length given by Ref. Neronov and Vovk (2010), and will improve once the lower bounds on these quantities becomes stricter, or if conclusive measurements of these quantities are made.
Acknowledgements
@@
This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. We acknowledge use of numpy (Harris et al., 2020), scipy (Virtanen et al., 2020), matplotlib (Hunter, 2007), astropy (Robitaille et al., 2013; Price-Whelan et al., 2018), jupyter (Kluyver et al., 2016), pandas (McKinney, 2010), seaborn (Waskom, 2021), bilby Ashton et al. (2019), bilby_pipe Romero-Shaw et al. (2020), dynesty Speagle (2020), pesummary Hoy and Raymond (2021), and lalsuite (LIGO Scientific Collaboration et al., 2018) software packages.
Appendix A Appendix A: Calculation of phase at detector.
The total additional time that a charged massive graviton would take to propagate from the source to the detector can be calculated by plugging in $\delta t^{m}_{f_{e},f_{e^{\prime}}}$ and $\delta t^{q}_{f_{e},f_{e^{\prime}}}$ of Eq.(8) and Eq.(9) in Eq.(7) resulting in
$\Delta t_{o}$ $\approx\Delta t_{e}+\frac{D_{L}c^{3}m_{g}^{2}}{2h^{2}}\left(\frac{1}{f_{e}^{2}}\right)$ (17)
 $\text{\hskip 71.13188pt}+\frac{2}{3}\frac{cD_{L}\lambda_{B}^{2}B^{2}q_{g}^{2}}{4!h^{2}}\left(\frac{1}{f_{e}^{2}}\right)-L(f_{e^{\prime}})$
where,
$L(f_{e^{\prime}})=\frac{D_{L}c^{3}m_{g}^{2}}{2h^{2}}\left(\frac{1}{f_{e^{\prime}}^{2}}\right)+\frac{2}{3}\frac{cD_{L}\lambda_{B}^{2}B^{2}q_{g}^{2}}{4!h^{2}}\left(\frac{1}{f_{e^{\prime}}^{2}}\right)\leavevmode\nobreak\ .$
Next, we may observe that the GW phasing in the frequency domain at the detector frame can be expressed as the following integral Blanchet et al. (2002); Blanchet (2014)
[…]
The contribution of the charge and the mass of a graviton on the GW phase can be obtained by replacing $t-t_{c}$ with $\Delta t_{o}$ from Eq. (17) and carrying out the integral. Under the assumption that the effect of the propagation over the large distance $D_{L}$ will dominate the corrections due to the back reaction in the modified theory Will (1998); Ghosh et al. (2023), we will get
$\psi(f)=$ $2\pi f\tilde{t}_{c}-\tilde{\Phi}_{c}-\frac{\pi}{4}+\sum_{n}\alpha^{n}f^{\frac{2n}{3}}$ (19)
 $\text{\hskip 42.67912pt}-\left\{\frac{\pi m_{g}^{2}c^{3}}{h^{2}}+\left(\frac{2}{3}B^{2}\lambda_{B}^{2}\right)\frac{2\pi q_{g}^{2}c}{4!h^{2}}\right\}\frac{D_{L}}{f}.$

### v3

@@ 3 Detector dependent phase shifts due to graviton propagation
$\Delta\phi=\frac{q_{g}\Phi}{\hbar c}~{}.$ (3.1)
Where $\Phi$ is the total flux in the region enclosed by the two trajectories. For randomly oriented magnetic fields with a typical magnitude $B$ and correlation length $\lambda_{B}$, we have $D_{L}/\lambda_{B}$ magnetic field domains between the source located at a distance $D_{L}$ away and the observer. Since the magnetic field domains are randomly oriented, only one-third of all the domains will contribute to the phase, behaving like a random walk. We can compute the mean distance from the origin for this random walk after $D_{L}/3\lambda_{B}$ steps to be $\sqrt{2D_{L}/3\pi\lambda_{B}}$. Then, the total flux in the triangular region as illustrated in Fig. (2) reads [17, 18]
$\Phi=\sqrt{\frac{D_{L}\lambda_{B}}{6\pi}}Bd~{},~{}q_{g}=\dfrac{\sqrt{6\pi}\hbar c}{\sqrt{\lambda_{B}}B}\left(\dfrac{\Delta\phi}{\sqrt{D_{L}}d}\right)$ (3.2)
@@
Here, $d$ is the separation between two detectors. The phase difference derived above is associated with distinct but coherent gravitons generated at the source observed at two different detectors. Nevertheless, similar to the case of photons, we can equate the classical GW phase difference observed and quantum phases of a graviton through the correspondence principle [17, 18]. So, by measuring the phase difference between the GW perturbations received at two detectors and associating this phase shift with the quantum phase shift of the graviton, we should be able to place bounds on the charge of the graviton using the above equation. Note that unlike the prescription described in the previous section, these bounds are free of any degeneracies with the mass of the graviton.
@@
In general, measuring $\Delta\phi$ between two detectors for a particular GW signal is difficult. For equal-mass, face-on GW signals that are well-described by the dominant $(\ell=m=2)$ harmonic of the GW signal, $\Delta\phi$ is perfectly degenerate with the reference phase at coalescence $\phi_{c}$, and any finite $\Delta\phi$ can be absorbed as a redefinition of $\phi_{c}$. However, when non-quadropolar modes are relevant for describing the GW signal (e.g. for unequal mass binaries or highly inclined binaries), the degeneracy between $\phi_{c}$ and $\Delta\phi$ is broken, and it is possible to measure $\Delta\phi$. This effect is already well-known in the literature in works dealing with strong gravitational lensing of GWs. Strong lensing naturally produces “Type-II” images that have a phase shift of $\pi/2$ relative to an un-lensed signal [28, 29], and this phase shift has been shown to be measurable in the presence of higher harmonics of the GW radiation, orbital precession and eccentricity in the GW signal [30, 31, 32, 33]. Hence, by making appropriate choices of the GW source, we can bound $\Delta\phi$ and thus the graviton charge independent of the mass.
4 Constraining the graviton charge
4.1 Assumptions on $B$ and $\lambda_{B}$
@@
It is apparent from Eq. (2.11) and Eq. (3.2) that the bounds on $q_{g}$ will become tighter with increasing $B$ and $\lambda_{B}$. While the value of $B$ and $\lambda_{B}$ is not known, many observations have placed lower limits on these parameters (see [34, 35] for a review). For instance, Ref. [36] obtained a lower limit on $B$ exploiting the non-observation of GeV gamma-ray emission from the electromagnetic cascade initiated by TeV gamma-rays (from blazars) to be $B>3\times 10^{-16}\,{\rm G}$ while being coherent over Mpc scales [36]. More specifically, these constraints state,
$B>\begin{cases}3\times 10^{-16}\left(\dfrac{\lambda_{B}}{1\,{\rm Mpc}}\right)^{-1/2}\,{\rm G}\,,&\lambda_{B}\leq 1\,{\rm Mpc}\\ 3\times 10^{-16}\,{\rm G}\,,&\lambda_{B}>1\,{\rm Mpc}\end{cases}$ (4.1)
@@
While there have been some improvements to these constraints over the years, we shall make a conservative choice $B=3\times 10^{-16}\,{\rm G}$ and $\lambda_{B}=1\,{\rm Mpc}$ for our results.
4.2 Results
@@
Figure 3: (Left) Joint posterior on $m_{g}$ and $q_{g}/e$ parameter space as inferred from GWTC-3 data. The colour corresponds to the probability density—darker colour signifies a higher probability density. Since the measurement of $q_{g}$ and $m_{g}$ are degenerate following Eq. (2.10), we see that the isoprobability contours are ellipses in this two-dimensional $m_{g}$$q_{g}/e$ plane. The black dashed line is the 90% isoprobability contour. Marginalizing over $m_{g}$, we infer a 90% upper limit of $|q_{g}|/e<3\times 10^{-34}{}$. (Right) Posterior probability distribution on $q_{g}/e$ as estimated from the measured phase difference between LIGO-Hanford and LIGO-Livingston for GW190814. At 90% CL, we place an upper limit of $2\times 10^{-26}{}$ on $q_{g}/e$. Although weaker than the constraints obtained from reinterpreting the graviton mass constraints, these constraints are independent and are not degenerate with constraints on the graviton mass.
@@
We use the posterior on the mass of the (uncharged) graviton combined over GWTC-3 events by LVK collaboration [12] and reinterpret it in the context of a charged massive graviton as discussed above. We report constraints on the joint $m_{g}$$q_{g}$ parameter space, with results shown in Fig. 3. As we see, the isoprobability contours in the image form ellipses, which follows directly from Eq. (2.11). We also show the ellipse corresponding to the $90\%$ contour. Under the aforementioned assumptions on $B$ and $\lambda_{B}$, the charge of the graviton is constrained to be
|$\absolutevalue{q_{g}}/e<3\times 10^{-34}{}\mbox{\quad.\quad}$ (4.2)
@@
We will next constrain the charge independently of the mass using Eq. (3.2). This method obviously has the advantage that we make no reference to the constraints on the graviton mass since the Aharonov-Bohm type phase shift can only occur in the presence of a non-zero charge. Even though this method is independent, the constraints are in general weaker than the one obtained using GW dephasing. As mentioned earlier, the measurement of phase difference between GW detectors can only be achieved with events that have significant higher harmonic content, failing which the phase difference measurement is exactly degenerate with the measurement of the coalescence phase of the binary $\phi_{c}$. We hence place bounds on $q_{g}$ using GW190814 [37], an asymmetric mass binary that showed significant evidence for higher-mode content in the signal. Although GW190814 was detected as a coincident signal by LIGO-Hanford, LIGO-Livingston, and Virgo detectors, we only use data from the LIGO detectors for our constraint due to the relatively low SNR in the Virgo detector222 While we have used only the phase difference between two detectors for our constraints with GW190814, this prescription can be easily generalized to multiple detectors by parametrizing all the relative phase differences through $q_{g}$. . We estimate the posterior on $\Delta\phi$ using the bilby [38] software package, using the nested sampler dynesty [39] to sample the posterior distribution. We use the waveform approximant IMRPhenomXPHM [40] implemented in lalsuite [41], and use the same standard priors as used in Ref. [37] along with a flat prior on $\Delta\phi$ in the range $[-\pi,\pi]$. We use the relative binning likelihood [42, 43, 44] implemented in bilby [45] to speed up the parameter inference [^3] While the relative binning implementation in bilby is strictly optimal for signals dominated by the quadrupolar mode of GW radiation, we have verified that the likelihood errors introduced by the non-optimality are small for our purposes. This is also in agreement with results of Ref. [45]., and use bilby_pipe [46] to streamline the inference runs. We obtain $\Delta\phi=-0.05^{+0.2}_{-0.25}$ at 90% CL which we consequently use to place an bound
|$\absolutevalue{q_{g}}/e<2\times 10^{-26}{}$ (4.3)
@@ 5 Conclusion
In this paper, we considered the possibility of a graviton endowed with a non-zero $U(1)$ charge. In the light of the Weinberg-Witten theorem, a non-zero $U(1)$ charge may make sense if the graviton is massive. A non-zero magnetic field will affect the trajectory of a charged graviton; we use this fact to compute the correction to the GW phasing that arises due to a non-zero charge. Our analysis points at a degeneracy of the graviton mass bounds from [12] with the possible graviton charge. Using the LVK bounds on the mass of the graviton, we can impose an upper bound on the U(1) charge of a graviton in terms of the electron charge as $|q_{g}|/e<3\times 10^{-34}{}$.
@@
In addition to the above bound, we also use the total phase difference between the GW detections at LIGO-Hanford and LIGO-Livingston for the event GW190814 to find another independent upper bound of the graviton charge. This is done by identifying the phase difference with the possible Aharonov-Bohm phase shift experienced by a $U(1)$ charged graviton [18, 17]. This provides us a novel bound on the graviton charge independent of its mass as $|q_{g}|/e<2\times 10^{-26}{}$. Similar to the bound obtained on the charge of a photon, the derivation of this independent bound on the graviton charge assumes the quantum phase difference arising due to a non-zero charge coincides with the classical phase difference of the two separate GW detections through the correspondence principle.
@@
In the near future, the constraints on $q_{g}$ will improve thanks to the larger baseline $d$ afforded by the addition of LIGO-India [47, 48] to the GW detector network. These constraints will also improve by a few orders of magnitude with even larger baselines and/or sensitivities afforded by the next generation of ground-based (Cosmic Explorer [49], Einstein Telescope [50], etc.) and space-based (LISA [51], DECIGO [52], etc.) detectors. Additionally, detections of more binaries with significant higher harmonics, eccentricity, and orbital precession will help in the measurement of the relative phase difference between detectors, hence also providing better constraints on $q_{g}$. Our analysis uses conservative estimates for the intergalactic magnetic field strength and the correlation length given by Ref. [36], and will improve once the lower bounds on these quantities becomes stricter, or if conclusive measurements of these quantities are made.
Appendix A Calculation of phase at detector.
The total additional time that a charged massive graviton would take to propagate from the source to the detector can be calculated by plugging in $\delta t^{m}_{f_{e},f_{e^{\prime}}}$ and $\delta t^{q}_{f_{e},f_{e^{\prime}}}$ of Eq.(2.8) and Eq.(2.9) in Eq.(2.7) resulting in
$\Delta t_{o}$ $\approx\Delta t_{e}+\frac{D_{L}c^{3}m_{g}^{2}}{2h^{2}}\left(\frac{1}{f_{e}^{2}}\right)+\frac{2}{3}\frac{cD_{L}\lambda_{B}^{2}B^{2}q_{g}^{2}}{4!h^{2}}\left(\frac{1}{f_{e}^{2}}\right)-L(f_{e^{\prime}})$ (A.1)
where,
$L(f_{e^{\prime}})=\frac{D_{L}c^{3}m_{g}^{2}}{2h^{2}}\left(\frac{1}{f_{e^{\prime}}^{2}}\right)+\frac{2}{3}\frac{cD_{L}\lambda_{B}^{2}B^{2}q_{g}^{2}}{4!h^{2}}\left(\frac{1}{f_{e^{\prime}}^{2}}\right)~{}.$
Next, we may observe that the GW phasing in the frequency domain at the detector frame can be expressed as the following integral [25, 26]
[…]
The contribution of the charge and the mass of a graviton on the GW phase can be obtained by replacing $t-t_{c}$ with $\Delta t_{o}$ from Eq. (A.1) and carrying out the integral. Under the assumption that the effect of the propagation over the large distance $D_{L}$ will dominate the corrections due to the back reaction in the modified theory [13, 27], we will get
$\psi(f)=$ $2\pi f\tilde{t}_{c}-\tilde{\Phi}_{c}-\frac{\pi}{4}+\sum_{n}\alpha^{n}f^{\frac{2n}{3}}-\left\{\frac{\pi m_{g}^{2}c^{3}}{h^{2}}+\left(\frac{2}{3}B^{2}\lambda_{B}^{2}\right)\frac{2\pi q_{g}^{2}c}{4!h^{2}}\right\}\frac{D_{L}}{f}.$ (A.3)
Acknowledgments
@@
This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. We acknowledge use of numpy [53], scipy [54], matplotlib [55], astropy [56, 57], jupyter [58], pandas [59], seaborn [60], bilby [38], bilby_pipe [46], dynesty [39], pesummary [61], and lalsuite [41] software packages.